# Boosting In-Context Learning in LLMs Through the Lens of Classical Supervised Learning

### v1

@@ 5 Conclusion
In this paper, we introduce Supervised Calibration (SC), a novel loss-minimization-based calibration framework designed to improve the performance of LLMs in ICL. We design SC to learn a class-specific affine transformation in logit space, allowing it to both shift and reorient the LLM’s decision boundary. Thanks to its expressive functional form, we show that SC generalizes and extends the corrective capabilities of many existing calibration methods for ICL. We further analyze SC through the lens of statistical learning theory, highlighting its approximation-estimation trade-off and providing theoretical insights into its strengths and limitations relative to prior approaches. Looking ahead, several avenues warrant exploration. Firstly, while our current ensembling strategy averages predictions over randomly selected contexts for each $i$-shot learner, future work could develop more principled approaches to context selection and weighting, which hold significant potential for improving SC’s performance. Secondly, although we provide theoretical intuition, a more rigorous theoretical analysis of SC, particularly accounting for the statistical dependencies introduced by our surrogate data generation, would be beneficial. Finally, while SC was developed for classification, extending its principles to calibrate LLMs for regression tasks presents an interesting and valuable direction for future research.
@@ Appendix A Implementation details
All large language models (LLMs) used in our experiments are based on publicly available implementations from the Hugging Face Transformers library [36]. We conduct all experiments on a dedicated computing node equipped with 8 NVIDIA A6000 Ada Generation GPUs.
Contextual Calibration [42](CC)
@@
Domain-Context Calibration [5]
We reproduce the DC baseline by using the test set as the unlabeled corpus to construct a bag-of-words. From this bag, we randomly sample tokens to create content-free and in-domain inputs with an average target length. Following [5], this process is repeated 20 times, and we compute the mean probability over these samples. Following the original implementation, we use this mean to normalize the LLM’s label-space probabilities computed for the test query and context.
Batch Calibration [44] (BC)
@@
We adopt an ensembling strategy for SC as outlined in Algorithm 2. For each configuration—$k=4$, $k=8$, and $k=16$—we set the minimum context size $i_{\text{min}}$ (as defined in Algorithm 2) to 1, and the maximum context size $i_{\text{max}}$ to $\min(5,k-1)$. We fix the regularization parameter $\lambda_{\text{inv}}$ to 10 across all settings and LLMs. Additionally, the number of context to be sampled from $\mathcal{C}(i)$ (given in Definition 1) for size $i$ during the inference is set as:
$m_{i}=\min\left(\left\lfloor\frac{|\mathcal{T}_{i}|}{2}\right\rfloor,24\right),$
where $|\mathcal{T}_{i}|$ denotes the number of available samples for context size $i$ outputted by Algorithm 1.
@@
$\tau=\operatorname{arccos}(\theta)$
We first compute the in-sample accuracy of the LLM while generating the training data through Algorithm 1. Based on this accuracy, we set the value of $\theta$ as follows:
$\theta=\begin{cases}\left(20^{\circ}\right)^{\frac{1}{K-1}}&\text{if accuracy}\geq 0.9\\ \left(45^{\circ}\right)^{\frac{1}{K-1}}&\text{if }0.7\leq\text{accuracy}<0.9\\ \left(90^{\circ}\right)^{\frac{1}{K-1}}&\text{if }0.5\leq\text{accuracy}<0.7\\ 180^{\circ}&\text{if accuracy}<0.5\end{cases}$
@@ Appendix B Additional related work
Mechanisms and prompt Optimization for ICL
Another line of work diagnoses why LLMs succeed or fail at ICL. The performance of a fixed prompt can swing from near random-guess to state of the art when the order of demonstrations is permuted [15], and it correlates strongly with the pre-training statistics of the tokens that appear in the prompt [22, 26]. From a theoretical perspective, ICL has been interpreted as implicit Bayesian inference in sequence models [37], while empirical evidence shows that sufficiently large models can even override entrenched semantic priors to learn arbitrary input–label mappings on the fly [35]. A complementary literature focuses on controlling these factors. Template-search methods [29, 19, 40] and example-selection algorithms [23, 14, 32] systematically pick demonstrations that maximize mutual information or diversity, while Wan and colleagues [31] add consistency and repetition checks. To make ICL more robust, researchers have proposed noisy–channel prompting [25], flipped learning that trains the model against label noise [39], k-nearest-neighbour label assignment [11], and lightweight decoder networks that adapt the prompt at inference time [4]. Together, these studies paint a converging picture: effective ICL hinges on matching the prompt (template and examples) to the model’s pre-training biases—then compensating for the remaining mismatches with task-specific selection or robust inference techniques.
@@ Appendix D Additional notation and detailed formulation
Definition 1 (Set of Ordered Contexts)
@@
Definition 2 (Set of Contexts Used for Query $x$)
@@
Definition 3 (Context Invariance Regularization Penalty)
@@
Table 2: Summary of Label Marginal based calibration methods. Each method adjusts the LLM prediction $P_{\text{LLM}}(y\mid x,C_{k})$ via the different estimators of $P_{\text{LLM}}(y|C_{k})$.
@@
Base LLM $\arg\max_{y}P_{\text{LLM}}(y\mid x,C_{k})$ Select the label with the highest conditional probability from the LLM.
Contextual Calibration (CC) $\arg\max_{y}\frac{P_{\text{LLM}}(y\mid x,C_{k})}{P_{\text{LLM}}(y\mid\text{NA},C_{k})}$ Normalize the prediction using a content-free input.
Domain-Context Calibration (DC) $\arg\max_{y}\frac{P_{\text{LLM}}(y\mid x,C_{k})}{\frac{1}{N}\sum_{i}P_{\text{LLM}}(y\mid\text{RandDom}_{i},C_{k})}$ Use randomly sampled domain prompts as a reference for normalization.
Batch Calibration (BC) $\arg\max_{y}\frac{P_{\text{LLM}}(y\mid x,C_{k})}{\frac{1}{N}\sum_{i}P_{\text{LLM}}(y\mid x_{i},C_{k})}$ Calibrate by averaging predictions over a batch of reference inputs.
Appendix E Surrogate data generation algorithm
@@
1:Demonstration set $C_{k}=\{(x^{(l)},y^{(l)})\}_{l=1}^{k}$ of size $k$.
2:Target context size $i$ such that $1\leq i<k$.
3:LM inference function $\text{Infer}(x,C_{i})$ that returns logit vector $\mathbf{m}(x;C_{i})$.
4:Initialize training set $\mathcal{T}_{i}\leftarrow\emptyset$.
5:Generate $\mathcal{C}(i)$, the set of all distinct ordered subsets of $C_{k}$ with size $i$. $\triangleright$ E.g., permutations of $C_{k}$, taking first $i$.
6:for each context $C^{(a)}_{i}\in\mathcal{C}(i)$ do
7:    Define the held-out set $R^{(a)}_{i}\leftarrow C_{k}\setminus C_{i}^{(a)}$. $\triangleright$ Set difference based on elements.
8:    for each query $(x,y)$ in $R^{(a)}_{i}$ do
9:        Compute model logits vector: $\mathbf{m}(x;C^{(a)}_{i})\leftarrow\text{Infer}(x,C^{(a)}_{i})$.
10:        Add to training set: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}\cup\{(m(x;C^{(a)}_{i}),y)\}$. $\triangleright$ Store feature vector and true label.
11:    end for
12:end for
13:Output: Training set $\mathcal{T}_{i}$ consisting of pairs (model logits, true label).
Appendix F Full algorithm
@@
1:Full demonstration set $C_{k}=\{(x^{(l)},y^{(l)})\}_{l=1}^{k}$; Set of context sizes $I=\{i_{min},\dots,i_{max}\}$; Regularization $\lambda_{inv}\geq 0$, $\tau\in[0,1]$; Context samples $m_{i}\geq 1$; Query $x$; Inference function $\text{Infer}(x,C)$ returns logit vector $\mathbf{m}(x,C)$.
2:Part 1: Training Phase
3:Initialize parameter set $\Theta\leftarrow\emptyset$.
4:for each context size $i\in I$ do
5:     Generate training data $\mathcal{T}_{i}$ using Algorithm 1 with $C_{k}$.
6:     Learn parameters $\hat{\bm{\theta}}^{i}$ by solving Eq. (3.3) using $\mathcal{T}_{i}$, $\lambda_{inv}$, $\tau$.
7:     Store $\hat{\bm{\theta}}^{i}$ in $\Theta$.
8:end for
9:Part 2: Prediction Phase (for query $x$)
10:Initialize list $P_{\text{list}}\leftarrow[]$.
11:for each context size $i\in I$ do
12:     Sample index set $\mathcal{M}_{i}\subseteq\{1,\dots,|\mathcal{C}(i)|\}$ uniformly at random such that $|\mathcal{M}_{i}|=m_{i}$.
13:     Retrieve learned parameters $\hat{\bm{\theta}}^{i}$ from $\Theta$.
14:     Retrieve sub-contexts $\{C_{i}^{(j)}\}_{j\in\mathcal{M}_{i}}$ from $\mathcal{C}(i)$ using $\mathcal{M}_{i}$.
15:     Initialize list $p_{\text{list}}^{(i)}\leftarrow[]$.
16:     for $j\in\mathcal{M}_{i}$ do
17:         $\mathbf{m}(x,C_{i}^{(j)})\leftarrow\text{Infer}(x,C_{i}^{(j)})$.
18:         $\mathbf{p}^{(j)}(x)\leftarrow\bm{f}(m(x,C_{i}^{(j)});\hat{\bm{\theta}}^{i})$.
19:         Append $\mathbf{p}^{(j)}(x)$ to $p_{\text{list}}^{(i)}$.
20:     end for
21:     $\hat{\mathbf{p}}_{i}(x)\leftarrow\frac{1}{m_{i}}\sum_{\mathbf{p}(x)\in p_{\text{list}}^{(i)}}\mathbf{p}(x)$.
22:     Append $\hat{\mathbf{p}}_{i}(x)$ to $P_{\text{list}}$.
23:end for
24:$\hat{\mathbf{p}}_{\text{SC{}}}(x)\leftarrow\frac{1}{|I|}\sum_{\mathbf{p}(x)\in P_{\text{list}}}\mathbf{p}(x)$.
25:Output: $\hat{y}_{\text{SC{}}}\in\arg\max_{y_{c}\in\mathcal{Y}}[\hat{\mathbf{p}}_{\text{SC{}}}(x)]_{c}$.
Appendix G Detailed numerical results
@@
Table 3: Average Macro-F1 scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 4-shot setting ($k=4$) over five random seeds. Values are presented as $\text{mean}_{\textit{standard deviation}}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 4: Average Macro-F1 scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 8-shot setting ($k=8$) over five random seeds. Values are presented as $\text{mean}_{\textit{standard deviation}}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 5: Average Macro-F1 scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 16-shot setting ($k=16$) over five random seeds. Values are presented as $\text{mean}_{\textit{standard deviation}}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 6: Average Accuracy scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 4-shot setting ($k=4$) over five random seeds. Values are presented as $\text{mean}_{\textit{standard deviation}}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 7: Average Accuracy scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 8-shot setting ($k=8$) over five random seeds. Values are presented as $\text{mean}_{\textit{standard deviation}}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 8: Average Accuracy scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 16-shot setting ($k=16$) over five random seeds. Values are presented as $\text{mean}_{\textit{standard deviation}}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Appendix H Ablation results
@@
First, to isolate the impact of learning the per-class scaling factor $w_{c}$, which underpins SC’s ability to reorient decision boundaries, we compare the full SC model against two alternatives: a restricted variant, $\text{SC}^{*}$ (where $w_{c}$ is fixed to 1, thus only learning an optimal bias term), and other baseline calibration methods. Our experiments reveal that $\text{SC}^{*}$ surpasses these other baselines. This suggests that estimating an optimal bias under SC framework is more effective than methods employed by LM methods. More critically, the full SC model achieves higher performance than $\text{SC}^{\ast}$, suggesting that the flexibility to learn the scaling factor—and therefore to both shift and rescale the LLM’s logits—offers a further advantage.
Second, we investigate whether aggregating calibrators trained with different context sizes improves predictive performance. Concretely, we train a collection of models $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$, where each calibrator is fitted using training data with $i$ in-context examples. We then ensemble these context-size-specific calibrators and evaluate the impact of increasing the number of distinct $i$-shot learners in the ensemble (i.e., increasing $|I|$). Empirically, we observe a consistent and monotonic improvement in both Accuracy and Macro-F1 scores as $|I|$ grows. This suggests that calibrators exposed to heterogeneous amounts of contextual information offer complementary signals, enhancing the robustness and predictive accuracy of the final calibrated output. These findings highlight a promising direction: with sufficient computational resources, one could train and ensemble an even broader set of context-specific calibrators to capture a richer diversity of contextual patterns, potentially unlocking further performance gains.
Third, we investigate the impact of the number of sampled sub-contexts ($m_{i}$) used for prediction averaging within each context-size-specific calibrator during the ensembling phase. Our findings reveal that increasing $m_{i}$—that is, averaging predictions over a greater number of distinct sub-contexts of size $i$—generally enhances Macro-F1 scores. This suggests that more comprehensive sampling of available context variations for each $i$-shot learner improves the accuracy of the ensemble’s output, helping to further reduce ICL’s sensitivity to specific context compositions.
@@
Figure 5: Accuracy and Macro-F1 scores of six methods on the Subjective dataset using the Mistral-7B-Instruct-v0.3 model in a 4-shot setting. Results are averaged over 5 random seeds. Bars represent the mean performance for each metric as indicated in the legend. $\text{SC}^{*}$ stands for the case where the scaling factor $w_{c}$ is fixed to 1 under the SC framework. Notably, SC consistently outperforms all other methods. The improved performance of $\text{SC}^{*}$ over other baselines suggests that estimating an optimal bias under SC framework is more effective than the methods employed by LM approaches, while the full SC further demonstrates the advantage of also learning the scaling factor.
Figure 6: Accuracy and Macro-F1 scores of six methods on the Subjective dataset using the Mistral-7B-Instruct-v0.3 model in an 8-shot setting. Results are averaged over 5 random seeds. Bars represent the mean performance for each metric as indicated in the legend. $\text{SC}^{*}$ stands for the case where the scaling factor $w_{c}$ is fixed to 1 under the SC framework. Notably, SC consistently outperforms all other methods. The improved performance of $\text{SC}^{*}$ over other baselines suggests that estimating an optimal bias under SC framework is more effective than the methods employed by LM approaches, while the full SC further demonstrates the advantage of also learning the scaling factor.
Figure 7: Impact of ensembling context-size-specific models within the SC framework on the Subjective dataset in an 8-shot setting. Result is reported for Mistral-7B-Instruct-v0.3, using Accuracy and Macro-F1 scores averaged over 5 random seeds. Each ensemble, denoted SC-N, aggregates calibration models trained on context sizes ranging from 1 to $N$ (e.g., SC-2 uses models with context sizes 1 and 2, SC-6 includes context sizes 1 through 6). The consistent improvement in performance as N increases across all three LLMs highlights the general benefit of aggregating insights from a more diverse set of k-shot learners.
@@
Figure 8: Impact of ensembling context-size-specific models within the SC framework on the Subjective dataset in an 8-shot setting. Results is reported for Qwen2-7B-Instruct, using Accuracy and Macro-F1 scores averaged over 5 random seeds. Each ensemble, denoted SC-N, aggregates calibration models trained on context sizes ranging from 1 to $N$ (e.g., SC-2 uses models with context sizes 1 and 2, SC-6 includes context sizes 1 through 6). The consistent improvement in performance as N increases across all three LLMs highlights the general benefit of aggregating insights from a more diverse set of k-shot learners.
@@
Figure 9: Impact of ensembling context-size-specific models within the SC framework on the Subjective dataset in an 8-shot setting. Result is reported for Llama-2-7b-chat-hf, using Accuracy and Macro-F1 scores averaged over 5 random seeds. Each ensemble, denoted SC-N, aggregates calibration models trained on context sizes ranging from 1 to $N$ (e.g., SC-2 uses models with context sizes 1 and 2, SC-6 includes context sizes 1 through 6). The consistent improvement in performance as N increases across all three LLMs highlights the general benefit of aggregating insights from a more diverse set of k-shot learners.
Figure 10: Impact of ensembling context-size-specific models within the SC framework on the Subjective dataset in an 16-shot setting. Result is reported for Llama-2-7b-chat-hf, using Accuracy and Macro-F1 scores averaged over 5 random seeds. Each ensemble, denoted SC-N, aggregates calibration models trained on context sizes ranging from 1 to $N$ (e.g., SC-2 uses models with context sizes 1 and 2, SC-11 includes context sizes 1 through 11). The consistent improvement in performance as N increases across all three LLMs highlights the general benefit of aggregating insights from a more diverse set of k-shot learners.

### v3

@@ 5 Conclusion
In this paper, we introduce Supervised Calibration (SC), a novel loss-minimization-based calibration framework designed to improve the performance of LLMs in ICL. We design SC to learn a class-specific affine transformation in logit space, allowing it to both shift and reorient the LLM’s decision boundary. Thanks to its expressive functional form, we show that SC generalizes and extends the corrective capabilities of many existing calibration methods for ICL. Looking ahead, several avenues warrant exploration. First, performance could be improved by developing more principled approaches to context selection and weighting, moving beyond the current random sampling strategy. Second, a more rigorous theoretical analysis of SC is needed, particularly one that accounts for the statistical dependencies introduced by our surrogate data generation method. Finally, extending the principles of SC to calibrate LLMs for regression tasks presents a valuable direction for future research.
Impact Statement
This paper introduces Supervised Calibration (SC), a calibration framework that improves the reliability of large language models used as few-shot classifiers in in-context learning. By learning a lightweight transformation of the model’s output scores using only the labeled examples already provided in the prompt, SC can reduce systematic prediction biases and make performance less brittle to context composition and ordering. In practical settings where few-shot classification is attractive because collecting large labeled datasets is costly, more stable and better-calibrated predictions can support safer and more dependable downstream pipelines.
At the same time, SC is not a substitute for broader safety and responsible-use practices. Calibrating a model can increase the confidence and apparent consistency of predictions, which may encourage over-reliance if deployed in high-stakes decision-making without appropriate oversight. Moreover, if the in-context labels are biased or unrepresentative, SC may reinforce those biases rather than remove them. We therefore recommend using SC alongside careful dataset and prompt curation, reporting performance variability across contexts, and maintaining human review or fallback mechanisms in sensitive applications. Finally, some SC configurations increase inference-time compute because they average predictions over multiple context variants; practitioners should weigh these accuracy gains against energy considerations.
@@ Appendix A Implementation Details
All large language models (LLMs) used in our experiments are based on publicly available implementations from the Hugging Face Transformers library (Wolf et al., 2020). We conduct all experiments on a dedicated computing node equipped with 8 NVIDIA A6000 Ada Generation GPUs.
Contextual Calibration (Zhao et al., 2021)(CC)
@@
Domain-Context Calibration (Fei et al., 2023)
We reproduce the DC baseline by using the test set as the unlabeled corpus to construct a bag-of-words. From this bag, we randomly sample tokens to create content-free and in-domain inputs with an average target length. This process is repeated 20 times, and we compute the mean probability over these samples. Following the original implementation, we use this mean to normalize the LLM’s label-space probabilities computed for the test query and context.
Batch Calibration (Zhou et al., 2023) (BC)
@@
We adopt an ensembling strategy for SC as outlined in Algorithm 2. For each configuration—$k=4$, $k=8$, and k=16k=16—we set the minimum context size $i_{\text{min}}$ (as defined in Algorithm 2) to 1, and the maximum context size $i_{\text{max}}$ to $\min(5,k-1)$. We fix the regularization parameter $\lambda_{\text{inv}}$ to 10 across all settings and LLMs. Additionally, the number of contexts to be sampled from $\mathcal{C}(i)$ (given in Definition 1) for size $i$ during the prediction is set as:
$m_{i}=\min\left(\left\lfloor\frac{\mathcal{T}_{i}}{2}\right\rfloor,24\right),$
where $\mathcal{T}_{i}$ denotes the number of available samples for context size $i$.
@@
$\tau=\operatorname{cos}(\alpha^{\circ})$
We first compute the in-sample accuracy of the LLM while generating the training data through Algorithm 1. Based on this accuracy, we set the value of $\alpha^{\circ}$ as follows:
$\alpha^{\circ}=\begin{cases}20^{\frac{1}{K-1}}&\text{if accuracy}\geq 0.9\\ 45^{\frac{1}{K-1}}&\text{if }0.7\leq\text{accuracy}<0.9\\ 90^{\frac{1}{K-1}}&\text{if }0.5\leq\text{accuracy}<0.7\\ 180&\text{if accuracy}<0.5\end{cases}$
@@ Appendix B Additional Related Work
Calibration via centroids.
A parallel line of work mitigates in-context biases by replacing the standard decision rule with centroid-based classification. Han et al. (2022) proposed Prototypical Calibration, which models output probability vectors using Gaussian mixtures and assigns labels based on cluster likelihood, improving robustness to prompt variation and class imbalance. Similarly, Cho et al. (2024) introduced Hidden Calibration, which operates in the model’s latent space by computing class centroids over hidden states and classifying based on proximity. Although these methods show empirical performance gains, they rely on additional data beyond the in-context examples, which may not always be available or compatible with the ICL setting.
@@ Appendix D Additional Notation and Detailed Formulation
Definition 1 (Set of Ordered Contexts).
@@
Definition 2 (Set of Contexts Used for Query $x$).
@@
Definition 3 (Context Invariance Regularization Penalty).
@@
Table 2: Summary of Label Based calibration methods. Each method adjusts the LLM prediction $P_{\text{LLM}}(y\mid x,C_{k})$ via the different estimators of $P_{\text{LLM}}(y|C_{k})$.
@@
LLM (Prob) $\arg\max_{y}P_{\text{LLM}}(y\mid x,C_{k})$ Selects the label with the highest conditional probability from the LLM.
Contextual Calibration (CC) $\arg\max_{y}\frac{P_{\text{LLM}}(y\mid x,C_{k})}{P_{\text{LLM}}(y\mid\text{NA},C_{k})}$ Normalizes the prediction using a content-free input to reduce label bias.
Domain-Context Calibration (DC) $\arg\max_{y}\frac{P_{\text{LLM}}(y\mid x,C_{k})}{\frac{1}{N}\sum_{i}P_{\text{LLM}}(y\mid\text{RandDom}_{i},C_{k})}$ Uses randomly sampled domain prompts as a reference for normalization.
Batch Calibration (BC) $\arg\max_{y}\frac{P_{\text{LLM}}(y\mid x,C_{k})}{\frac{1}{N}\sum_{i}P_{\text{LLM}}(y\mid x_{i},C_{k})}$ Calibrates by averaging predictions over a batch of reference inputs.
Appendix E Full Algorithms
@@
  Input: Demonstration set $C_{k}=\{(x^{(l)},y^{(l)})\}_{l=1}^{k}$ of size $k$; target context size $i$ such that $1\leq i<k$; LM inference function $\text{Infer}(x,C)$ that returns logit vector $\mathbf{m}(x;C)$.
  Initialize training set $\mathcal{T}_{i}\leftarrow\emptyset$.
  Generate $\mathcal{C}(i)$, the set of all distinct ordered subsets of $C_{k}$ with size $i$ (e.g., permutations of $C_{k}$, taking first ii).
  for each context $C_{i}^{(a)}\in\mathcal{C}(i)$ do
  Define the held-out set $R_{i}^{(a)}\leftarrow C_{k}\setminus C_{i}^{(a)}$ (set difference based on elements).
  for each query $(x,y)$ in $R_{i}^{(a)}$ do
   Compute model logits: $\mathbf{m}(x;C_{i}^{(a)})\leftarrow\text{Infer}(x,C_{i}^{(a)})$.
   $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}\cup\{(\mathbf{m}(x;C_{i}^{(a)}),y)\}$ (store feature vector and true label).
  end for
  end for
  Output: Training set $\mathcal{T}_{i}$ consisting of pairs (model logits, true label).
@@
  Input: Full demonstration set $C_{k}=\{(x^{(l)},y^{(l)})\}_{l=1}^{k}$; context sizes $I=\{i_{\min},\dots,i_{\max}\}$; regularization $\lambda_{\mathrm{inv}}\geq 0$, $\tau\in[0,1]$; context samples $m_{i}\geq 1$; query $x$; inference function $\text{Infer}(x,C)$ returns logit vector $\mathbf{m}(x,C)$.
  Part 1: Training Phase
  Initialize parameter set $\Theta\leftarrow\emptyset$.
  for each context size $i\in I$ do
  Generate training data $\mathcal{T}_{i}$ using Algorithm 1 with $C_{k}$.
  Learn parameters $\hat{\bm{\theta}}^{i}$ by solving Eq. (3.3) using $\mathcal{T}_{i}$, $\lambda_{\mathrm{inv}}$, $\tau$.
  Store $\hat{\bm{\theta}}^{i}$ in $\Theta$.
  end for
  Part 2: Prediction Phase (for query $x$)
  Initialize list $P_{\text{list}}\leftarrow[\,]$.
  for each context size $i\in I$ do
  Sample index set $\mathcal{M}_{i}\subseteq\{1,\dots,|\mathcal{C}(i)|\}$ uniformly at random such that $|\mathcal{M}_{i}|=m_{i}$.
  Retrieve learned parameters $\hat{\bm{\theta}}^{i}$ from $\Theta$.
  Retrieve sub-contexts $\{C_{i}^{(j)}\}_{j\in\mathcal{M}_{i}}$ from $\mathcal{C}(i)$ using $\mathcal{M}_{i}$.
  Initialize list $p_{\text{list}}^{(i)}\leftarrow[\,]$.
  for each $j\in\mathcal{M}_{i}$ do
   $\mathbf{m}(x,C_{i}^{(j)})\leftarrow\text{Infer}(x,C_{i}^{(j)})$.
   $\mathbf{p}^{(j)}(x)\leftarrow\bm{f}(\mathbf{m}(x,C_{i}^{(j)});\hat{\bm{\theta}}^{i})$.
   Append $\mathbf{p}^{(j)}(x)$ to $p_{\text{list}}^{(i)}$.
  end for
  $\hat{\mathbf{p}}_{i}(x)\leftarrow\frac{1}{m_{i}}\sum_{\mathbf{p}(x)\in p_{\text{list}}^{(i)}}\mathbf{p}(x)$.
  Append $\hat{\mathbf{p}}_{i}(x)$ to $P_{\text{list}}$.
  end for
  $\hat{\mathbf{p}}_{\text{SC{}}}(x)\leftarrow\frac{1}{|I|}\sum_{\mathbf{p}(x)\in P_{\text{list}}}\mathbf{p}(x)$.
  Output: $\hat{y}_{\text{SC{}}}\in\arg\max_{y_{c}\in\mathcal{Y}}[\hat{\mathbf{p}}_{\text{SC{}}}(x)]_{c}$.
Appendix F Detailed Numerical Results
@@
Table 3: Average Macro-F1 scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 4-shot setting ($k=4$) over five random seeds. Values are presented as $\text{mean}_{s.d}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 4: Average Macro-F1 scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 8-shot setting ($k=8$) over five random seeds. Values are presented as $\text{mean}_{s.d}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 5: Average Macro-F1 scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 16-shot setting ($k=16$) over five random seeds. Values are presented as $\text{mean}_{s.d}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 6: Average Accuracy scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 4-shot setting ($k=4$) over five random seeds. Values are presented as $\text{mean}_{s.d}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 7: Average Accuracy scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 8-shot setting ($k=8$) over five random seeds. Values are presented as $\text{mean}_{s.d}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Table 8: Average Accuracy scores (%) for various calibration methods on selected datasets, evaluated for each LLM in the 16-shot setting ($k=16$) over five random seeds. Values are presented as $\text{mean}_{s.d}$, with the highest score in each column highlighted in bold and shaded gray.
@@
Appendix G Ablation Results
@@ G.1 Scaling Matters
First, to isolate the impact of learning the per-class scaling factor $w_{c}$, which underpins SC’s ability to reorient decision boundaries, we compare the full SC model against two alternatives in Figure 5: a restricted variant, $\text{SC}^{*}$ (where $w_{c}$ is fixed to 1, thus only learning an optimal bias term), and other baseline calibration methods. Our experiments reveal that $\text{SC}^{*}$ surpasses these other baselines. This suggests that estimating an optimal bias under SC framework is more effective than methods employed by LM methods. More critically, the full SC model achieves higher performance than $\text{SC}^{\ast}$, suggesting that the flexibility to learn the scaling factor—and therefore to both shift and rescale the LLM’s logits—offers a further advantage.
The performance difference between SC and $\text{SC}^{\ast}$ is particularly apparent on a challenging 8-shot, multi-class classification task (SST-5) where the base model’s predictions are often poorly oriented. Specifically, Table 9 shows that $\text{SC}^{\ast}$ method achieves a very low Macro-F1 of 0.1004, indicating its inability to correct the model’s predictions. In stark contrast, the full SC method boosts the Macro-F1 to 0.4106 and accuracy to 0.4352, representing a four-fold improvement. This vast performance gap confirms our hypothesis: on difficult tasks with severe miscalibrations, only full SC, capable of both shifting and scaling the decision boundary, can effectively correct severly misaligned LLM.
@@
Figure 5: Accuracy and Macro-F1 scores of six methods on the Subjective dataset using the Mistral-7B-Instruct-v0.3 model in (a) 4-shot and (b) 8-shot settings. Results are averaged over 5 random seeds. Bars represent the mean performance for each metric as indicated in the legend. $\text{SC}^{*}$ stands for the case where the scaling factor $w_{c}$ is fixed to 1 under the SC framework. Notably, SC consistently outperforms all other methods in both settings. The improved performance of $\text{SC}^{*}$ over other baselines suggests that estimating an optimal bias under SC framework is more effective than the methods employed by LM approaches, while the full SC further demonstrates the advantage of also learning the scaling factor.
Table 9: Comparison on the 8-shot SST-5 task with the Qwen2-7B-Instruct model. SC v.s SC*.
Method Macro-F1 (mean $\pm$ SE) Accuracy (mean $\pm$ SE)
Base LLM $0.1565\pm 0.0033$ $0.2434\pm 0.0016$
SC* (scaling=1) $0.1004\pm 0.0125$ $0.2227\pm 0.0168$
SC $0.4106\pm 0.0280$ $0.4352\pm 0.0428$
G.2 Ensembling Across Context Sizes ($|I|$) Improves Performance
Second, we investigate whether ensembling calibrators trained with different context sizes improves predictive performance. Concretely, we train a collection of models $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$, where each calibrator is fitted using training data with $i$ in-context examples. We then ensemble these context-size-specific calibrators and evaluate the impact of increasing the number of distinct $i$-shot learners in the ensemble (i.e., increasing $|I|$). Empirically, we observe a consistent and monotonic improvement in both Accuracy and Macro-F1 scores as $|I|$ grows as shown in Figure 6 and 7. This suggests that calibrators exposed to heterogeneous amounts of contextual information offer complementary signals, enhancing the robustness and predictive accuracy of the final calibrated output. These findings highlight a promising direction: with sufficient computational resources, one could train and ensemble an even broader set of context-specific calibrators to capture a richer diversity of contextual patterns, potentially unlocking further performance gains.
@@
Figure 6: Impact of ensembling context-size-specific models within the SC framework on the Subjective dataset in an 8-shot setting. Results are reported for (a) Llama-2-7b-chat-hf, (b) Qwen2-7B-Instruct, and (c) Mistral-7B-Instruct-v0.3, using Accuracy and Macro-F1 scores averaged over 5 random seeds. Each ensemble, denoted SC-N, aggregates calibration models trained on context sizes ranging from 1 to NN (e.g., SC-2 uses models with context sizes 1 and 2, SC-6 includes context sizes 1 through 6). The consistent improvement in performance as N increases across all three LLMs highlights the general benefit of aggregating insights from a more diverse set of k-shot learners.
@@
Figure 7: Impact of ensembling context-size-specific models within the SC framework on the Subjective dataset in a 16-shot setting. Result is reported for Llama-2-7b-chat-hf, using Accuracy and Macro-F1 scores averaged over 5 random seeds. Each ensemble, denoted SC-N, aggregates calibration models trained on context sizes ranging from 1 to NN (e.g., SC-2 uses models with context sizes 1 and 2, SC-11 includes context sizes 1 through 11). The consistent improvement in performance as N increases across all three LLMs highlights the general benefit of aggregating insights from a more diverse set of k-shot learners.
G.3 Macro-F1 Gains as the number of sampled sub-contexts Increases
Next, we investigate the impact of the number of sampled sub-contexts ($m_{i}$) used for prediction averaging within each context-size-specific calibrator during the ensembling phase. In Figure 8, our findings reveal that increasing $m_{i}$ (i.e., averaging predictions over a greater number of distinct sub-contexts of size $i$) generally enhances Macro-F1 scores. This suggests that more comprehensive sampling of available context variations for each $i$-shot learner improves the accuracy of the ensemble’s output, helping to further reduce ICL’s sensitivity to specific context compositions.